OpenStudy (anonymous):

Neophyte What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 20-in TV set also has a width of 16 inches. What is it's height in inches?

OpenStudy (saifoo.khan):

Use pythagoas!

OpenStudy (saifoo.khan):

Pythagoras*

OpenStudy (anonymous):

it's the same as the previous question we did

OpenStudy (saifoo.khan):

aaah!

OpenStudy (anonymous):

so it's still 30, right?

OpenStudy (saifoo.khan):

\[25^2−16^2=h^2\] h = 20

OpenStudy (saifoo.khan):

its 20.

OpenStudy (anonymous):

oh. Sorry guys. I have serious issues with my short term memory.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

I know. lol it sucks. I always have to try & relearn the steps.